OpenStudy (anonymous):

One leg of a right triangle has length 7. The lengths of the other two sides are whole numbers. The length of the other leg is___and the length of the hypotenuse is_______

OpenStudy (anonymous):

The length of the other leg is 24 and the length of the hypotenuse is 25.

OpenStudy (anonymous):

I don't think that answer is right..Js

OpenStudy (anonymous):

\[25^2=24^2+7^2 \]\[625=576+49 \]\[625=625 \]

OpenStudy (anonymous):

thank you!!!!